Find the value for \tan\:\theta tan θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Find the value for \sec\:\theta sec θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Find the value for \csc\:\theta csc θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Find the value for \tan\:\theta tan θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Find the value for \cos\:\theta cos θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Find the value for \sin\:\theta sin θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.

Answer by stanbon(75887)

Find the value for \tan\:\theta tan θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Since tan = y/x, y = -3 and x = 4
Then r = sqrt[(-3)^2+4^2] = sqrt(25) = 5
-------------------------------------------
Find the value for \sec\:\theta sec θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
Note:: sec = r/x
---------------------------
Find the value for \csc\:\theta csc θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form
Note: csc = r/y
--------------------------
Find the value for \tan\:\theta tan θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
tan = 4/(-3) so y = 4 and x = -3
And r = 5
Find the value for \cos\:\theta cos θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
cos = x/r
-----
Find the value for \sin\:\theta sin θ given the point (-3, 4) on the terminal side. Leave your answer in fraction form.
sin = y/r
----------------
Cheers,
Stan H.
-----------------